OpenStudy (anonymous):

If f(x)=x12−x and g(x)=2x3−x12−x, find f(x)−g(x). −2x3−2x−2x12 −2x3−2x 2x3−2x+2x12 −2x3+2x12

OpenStudy (anonymous):

what is f(x) ?

OpenStudy (haseeb96):

f(x)-g(x)= x12-x -(2x3-x12-x) =x12 -x -2x3 +x12 +x =-2x3 +2x12 is your answer

OpenStudy (anonymous):

thanks got it right will you help me with my next one?